# A Tutorial on MIMO-OFDM ISAC: From Far-Field to Near-Field

### v1

@@ IV-E1 Complexity Analysis
Figure 11: Estimation accuracy of the considered 1D algorithms in the three individual domains for MIMO-OFDM ISAC.
@@ IV-E2 Estimation Accuracy and Resolution Analysis
To study the relative performance of the 1D parameter estimation algorithms discussed above, we consider a scenario similar to that outlined in Table III, but with the spacing between the parameters increased to guarantee that all the algorithms resolve the targets. We conducted 300 Monte Carlo experiments to estimate parameters in three domains via these 1D parameter estimation algorithms for the SNR in (8) varying from -40 dB to 20 dB. The root mean-squared error (RMSE) for each algorithm in the three domains is shown in Fig. 11.
We observe a threshold effect at approximately -20 dB where the algorithm performance improves dramatically for the above setup. Among them, the periodogram shows the best performance, followed by MUSIC, with ESPRIT being the least estimation accuracy. The performance of PM and ESPRIT in the spatial domain surpasses that in the delay and Doppler domains since ICI affects the latter, while signals in spatial domain remain approximately orthogonal. For delay estimation, PM-MUSIC fails due to signal coherence at high SNR, although this can be addressed by spatial smoothing. FFT-MUSIC and MUSIC have similar estimation accuracy, while ROOT-MUSIC performs poorly at low SNR but approaches MUSIC at high SNR. Finally, TLS-ESPRIT has slightly higher accuracy than LS-ESPRIT at low SNR. PM-ESPRIT exhibits superior performance exclusively at high SNR, particularly for delay and Doppler estimation.
@@
Regarding resolution, we derived the theoretical resolution of the periodogram algorithm for AoA, delay, and Doppler estimation in (41) and (42). However, deriving the theoretical resolution for super-resolution algorithms such as MUSIC and ESPRIT is challenging. Therefore, we design a simulation experiment to study this issue. Using the estimation error of two closely positioned targets as a constraint, we statistically determine the probability of successfully resolving the two targets for each of the considered algorithm to analyze their resolution. Taking the angular resolution as an example, the Sen-RX is set as a ULA equipped with $M_{r}=16$ antennas with half-wavelength element-spacing. The ISAC-TX transmits an OFDM signal with $N=128$, the CPI $P=64$ symbols, subcarrier spacing $\Delta f=120\ \mathrm{kHz}$, and CP duration $T_{\mathrm{cp}}=\frac{1}{4}T$. Since the theoretical resolution of the periodogram is $7.1^{\circ}$ under these conditions, we establish two targets, for which the AoA of the first target is $0^{\circ}$, and that of the second one varies from $0.2^{\circ}$ to $9^{\circ}$. An algorithm is considered to have resolved the two targets successfully if the RMSE is within $0.5^{\circ}$ for $\mathrm{SNR}=-10\ \mathrm{dB}$, and within $0.3^{\circ}$ for $\mathrm{SNR}=10\ \mathrm{dB}$. As the angle of the second target changes, the probabilities of successfully resolving the two targets using the periodogram, MUSIC, and ESPRIT algorithms for $\mathrm{SNR}=-10\mathrm{dB}$ and $\mathrm{SNR}=10\mathrm{dB}$ are shown in Figs. 12(a) and 12(b), respectively, based on $200$ Monte Carlo experiments.
The simulation results in Figs. 12(a) and 12(b) show the angular resolution of the periodogram algorithm is $7^{\circ}$ for both $\mathrm{SNR}=-10\mathrm{dB}$ and $\mathrm{SNR}=10\mathrm{dB}$, consistent with the theoretical resolution in (41). The resolution of the MUSIC and ESPRIT algorithms depends on the SNR. For lower SNR, MUSIC and ESPRIT exhibit similar resolution, approximately $4^{\circ}$. For higher SNR, ESPRIT achieves superior resolution around $0.6^{\circ}$, whereas MUSIC provides resolution around $1^{\circ}$. Overall, compared to the periodogram, super-resolution algorithms significantly enhance parameter estimation resolution, particularly at high SNR levels.
@@ V Near-field MIMO-OFDM ISAC
In this section, we extend the results in Section IV to the case where the targets are located in the near-field region of the Sen-RX. While the main signal processing strategies outlined in Fig. 4 are still valid in the near-field, the spatial domain estimation must be extended from AoA estimation only in far-field to joint angle-range estimation in the near-field by exploiting the characteristics of spherical wavefronts. Therefore, this section focuses mainly on spatial domain estimation.
@@ V-A Near-field Sensing Signal Model
As shown in Fig. 13, we assume that the Sen-RX is equipped with an extremely large-scale ULA, which is placed along the $y$-axis with $\varepsilon_{r}=0,$ and $\varepsilon_{m}=m-(M_{r}+1)/2$ in (14). The corresponding steering matrix is denoted as $\mathbf{A}_{\rm{R}}(r,\theta)=[\mathbf{a}_{\mathrm{R}}(r_{1},\theta_{1}),...,\mathbf{a}_{\mathrm{R}}(r_{K},\theta_{K})]$. Based on the model in (22), the $M_{r}\times Q_{N,P}$ received signal is given by
@@
where $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{K}]^{T}\in\mathbb{C}^{K\times Q_{N,P}}$ denotes the equivalent reflected target signals described in (23). Note that compared with the far-field model in (22), the array response vector for near-field sensing is a function of both the target angle and distance, hence allowing simultaneous estimation of both parameters independent of delay-based distance estimation. This will not only improve the distance estimation accuracy (especially if the distance estimation is performed jointly in the AoA and delay domains), but it can also alleviate the synchronization errors that typically arise in traditional delay-based distance estimation techniques. In order to estimate the angle and distance parameters based on spatial-domain processing, several methods are introduced below, which can be generally categorized into directly applying far-field algorithms, extending far-field algorithms to three types, namely, near-field, and low-complexity near-field algorithms.
V-B Directly Applying Far-field Algorithms to Near-field Models
In this section, we illustrate the performance of directly applying the far-field algorithms discussed in Section IV-B while ignoring near-field effects. The result for the standard Periodogram is shown in Fig. 14, where $K=2$ targets are considered. We see that only the target with $r=150m$ has an accurate AoA estimate based on (45). For the closer target at $r=30m$, multiple peaks exist around the ground-truth AoA, which is due to the mismatch between the near-field channel and the assumed far-field model. This is known as the so-called energy spread effect, which illustrates the necessity of developing dedicated near-field target localization algorithms.
@@ V-B Directly Applying Far-field Algorithms to Near-field Models
V-C Extending Far-field Algorithms to the Near-field
@@ V-C1 Near-field Beam Focusing
For near-field models, the distance and angle parameters are closely coupled in the near-field array response vector. In addition, the non-linear phase variation from antenna to antenna in near-field models also means that the IFFT is no longer directly applicable. Therefore, we introduce 2D beamforming in both the angle and distance domains in this section. Based on (74), the spatial domain sample covariance matrix of $\bf X$ is given by
@@
The peaks of ${P_{\mathrm{2D-BF}}}$ in the 2-D angle-distance can be used to estimate the target angle and distance parameters.
@@ V-C2 Near-field MUSIC
Similar to (47), by performing EVD to the covariance matrix ${\bf R}_{X}$ and obtaining the noise subspace ${\bf E}_{n}$, the 2-D MUSIC spectrum can be expressed as
@@ V-C4 Simulations and discussions
Figs. 15(a) and 15(b) show the 2D spectra for near-field beam focusing and MUSIC respectively. The number of antennas is $M_{r}=256$, and the locations of the two targets are $(r_{1},\theta_{1})=(5\mathrm{m},10^{\circ})$ and $(r_{2},\theta_{2})=(10\mathrm{m},20^{\circ})$, respectively. It is observed that both beamforming and MUSIC can localize the closer target with higher accuracy than the distant target, primarily due to the limited accuracy with respect to distance.
@@
Figure 15: Near-field angle-distance estimation by extending far-field algorithms to near-field.
We see that several issues arise that complicate parameter estimation in near-field scenarios. Besides the fact that some algorithms can no longer be directly applied, we see that 2D rather than 1D searches are necessary to jointly estimate the AoAs and ranges, which adds complexity. The accuracy of the range estimate is considerably less reliable than the AoA estimate, especially as the range increases to the point where the far-field model holds [127]. This is due to the decreasing sensitivity of the model in such scenarios, but is a factor that must be taken into account when jointly estimating the range by including the delay domain.
To address these challenges, one feasible approach is to decouple estimation of the angle and distance parameters to reduce the complexity of implementing the parameter optimization. Based on this idea, the following low-complexity near-field sensing algorithms are presented.
@@ V-D Low-complexity Near-Field Estimation Algorithms
Recently, several efficient algorithms have been proposed to improve the accuracy and resolution of near-field sensing with low computational complexity by efficiently decoupling the distance and angle parameters. To that end, most existing works have considered the approximate Fresnel array response model in (16) based on second-order Taylor approximation, and the parameters to be estimated are transformed from $\theta_{k}$ and $r_{k}$ to ${\omega_{k}}=-\frac{{2\pi{d}\sin{\theta_{k}}}}{\lambda}$ and ${\psi_{k}}=\frac{{\pi d^{2}{{\cos}^{2}}{\theta_{k}}}}{{\lambda{r_{k}}}}$, respectively.
@@ V-D1 Second-order cumulant (SoC)
Note that in (16), $\omega_{k}$ only relates to $\theta_{k}$, while $\psi_{k}$ depends on both $\theta_{k}$ and $r_{k}$. Therefore, dedicated second-order culumants can be constructed to transform the multi-dimensional search into two single-parameter estimation problems. Based on (74) and (16), the noise-free correlation sequence between the $(am+b)$th and $m$th antenna elements can be written as [128]
@@
where ${{\bf{A}}_{R,m}}$ is the $m$th row of ${{\bf{A}}_{R}}$. Approximation $(a)$ holds by assuming ${\bf{s}}_{p}^{H}{{\bf{s}}_{q}}\approx 0,p\neq q$. $J=\frac{M_{r}-1}{2}$ and $p_{k}=\frac{1}{Q_{N,P}}{\bf{S}}(k,:){{\bf{S}}^{H}}(k,:)$ denotes the equivalent transmitted signal. Furthermore, by properly choosing $a$ and $b$, $\psi_{k}$ can be eliminated from the resulting expression, and estimation of the distance and angle parameters can be decoupled. For instance, substituting $(a,b)=(-1,0)$ and $(a,b)=(1,1)$ into (V-D1), we obtain the following correlation sequences
@@ V-D3 FFT-enhanced low-complexity and super-resolution algorithm
In order to further reduce the complexity, the authors in [133] proposed an efficient algorithm to narrow the search region for 2D angle-distance estimation in (77). The main idea of this algorithm is to first eliminate those regions that do not contain any targets. Afterwards, 2D MUSIC is performed within a region of reduced size in order to estimate the angle and distance, hence effectively balancing computational complexity and super-resolution. This algorithm involves three steps:
FFT-enhanced angle cluster determination: Applying an $S$-point FFT and IFFT to the columns and rows of sample covariance matrix ${\bf R}_{X}$ in (75), respectively, the output can be used to eliminate angle regions where no targets exist. The resulting spectrum can be expressed as ${{\bf{p}}}=\text{diag}({\bf{W}}{{\bf{R}}_{X}}{{\bf{W}}^{-1}}),$ where $\bf W$ is the DFT matrix, and diag(A) means extracting the diagonal elements of $\bf A$ to form a vector. The $n$th angle cluster is represented by the minimum and maximum angle value within this cluster, denoted as ${\bm{\alpha}}_{n}=[{\underline{\alpha}_{n}},{{\overline{\alpha}}_{n}}],n=1,...,L$. $L$ is the number of clusters.
Distance cluster determination: For each angle cluster ${\bm{\alpha}}_{n}=[{\underline{\alpha}_{n}},{{\overline{\alpha}}_{n}}]$, 1D beamforming is performed over the distance domain for $\underline{\alpha_{n}}$ and ${{\overline{\alpha}}_{n}}$, respectively.
@@
By setting a threshold $\Gamma$, the distance cluster ${{{\bm{\underline{\beta}}}}_{n}}$ can be expressed as $P_{low}(r)<\Gamma$, and ${{{\bm{\bar{\beta}}}}_{n}}$ can be obtained similarly. Therefore, the overall distance cluster is expressed as ${\bm{\beta}}_{n}={{{\bm{\underline{\beta}}}}_{n}}\cup{{{\bm{\bar{\beta}}}}_{n}}.$
Super-resolution 2D sensing: Finally, for each cluster $n$, 2D MUSIC is performed for $\theta\in{\bm{\alpha}}_{n}$ and $r\in{\bm{\beta}}_{n}$. By clustering the angles and distance, the number of grids for the 2D search decreases significantly, thus reducing the overall complexity.

### v2

@@ IV-E1 Complexity Analysis
Figure 16: Estimation accuracy of various 1D algorithms in the three individual domains for MIMO-OFDM ISAC.
As shown in Table VII, when the number of antennas is small and the number of subcarriers and OFDM symbols is large, numerous snapshots are available for AoA estimation, and the periodogram algorithm even exhibits higher complexity than MUSIC and ESPRIT. However, the situation changes as the number of antennas increases. The periodogram becomes more computationally efficient due to the fast computation of the FFT, followed by ESPRIT and MUSIC, while OMP has the highest complexity. The periodogram and OMP can estimate the desired parameters using only partial snapshots or even a single snapshot to reduce complexity, at the cost of increased error. Among the variants of MUSIC, FFT-MUSIC has the lowest complexity, followed by ROOT-MUSIC, with PM-MUSIC offering the least reduction in complexity. Total LS (TLS)-ESPRIT [80] is a variation of LS-ESPRIT that offers slightly improved performance, but with a corresponding slight increase in complexity. Unlike MUSIC, the PM algorithm has the potential to significantly reduce the complexity of ESPRIT. Overall, the periodogram, FFT-MUSIC, and PM-ESPRIT algorithms offer the lowest complexity.
The complexity of 2D algorithms for joint delay-Doppler estimation and 3D algorithms for joint AoA-delay-Doppler estimation is presented in Table VIII. Since the complexity analysis for the 2D/3D algorithms follows a similar approach to that of 1D algorithms, it will not be elaborated on further here. Note that 2D/3D algorithms not only have a high computational complexity but also involve calculations with large-dimensional matrices, imposing significant computational demands. To achieve high-accuracy parameter estimation while mitigating these challenges, coarse-fine grid search [138] and signal decimation [81, 139] techniques can significantly reduce both computational complexity and hardware requirements.
@@ IV-E2 Estimation Accuracy and Resolution Analysis
To study the performance of the 1D parameter estimation algorithms discussed above, we consider a scenario similar to that outlined in Table V, but with the spacing between the parameters increased to guarantee that all the algorithms resolve the targets. We conducted 300 Monte Carlo experiments to estimate the target parameters in three domains via these 1D parameter estimation algorithms for the SNR in (8) varying from -40 dB to 20 dB. The root mean-squared error (RMSE) for each algorithm in the three domains is shown in Fig. 16.
We observe a threshold effect at approximately -20 dB where the algorithm performance improves dramatically for the above setup. The periodogram shows the best performance, followed by MUSIC, with ESPRIT providing the lowest estimation accuracy. The performance of PM and ESPRIT in the spatial domain surpasses that in the delay and Doppler domains since ICI affects the latter, while signals in spatial domain remain approximately orthogonal. For delay estimation, PM-MUSIC fails due to signal coherence at high SNR, although this can be addressed by spatial smoothing. FFT-MUSIC and MUSIC have similar estimation accuracy, while ROOT-MUSIC performs poorly at low SNR but approaches MUSIC at high SNR. Finally, TLS-ESPRIT has slightly higher accuracy than LS-ESPRIT at low SNR. PM-ESPRIT exhibits superior performance exclusively at high SNR, particularly for delay and Doppler estimation.
Earlier, we obtained the theoretical resolution of the periodogram algorithm for AoA, delay, and Doppler estimation in (41) and (42). However, deriving the theoretical resolution for super-resolution algorithms such as MUSIC and ESPRIT is challenging. Therefore, we design a simulation experiment to study this issue. Using the estimation error of two closely positioned targets as a constraint, we statistically determine the probability of successfully resolving the two targets for each of the considered algorithms. Taking the angular resolution as an example, the Sen-RX is chosen to be a ULA with $M_{r}=16$ half-wavelength-spaced antennas. The ISAC-TX transmits an OFDM signal with $N=128$, the CPI is $P=64$ symbols, subcarrier spacing $\Delta f=120\ \mathrm{kHz}$, and CP duration $T_{\mathrm{cp}}=\frac{1}{4}T$. Since the theoretical resolution of the periodogram is $7.1^{\circ}$ under these conditions, we assume one target located at $0^{\circ}$ and the other at different angles between $0.2^{\circ}$ and $9^{\circ}$. An algorithm is considered to have resolved the two targets successfully if the RMSE is within $0.5^{\circ}$ for $\mathrm{SNR}=-10\ \mathrm{dB}$, and within $0.3^{\circ}$ for $\mathrm{SNR}=10\ \mathrm{dB}$. As the angle of the second target changes, the probabilities of successfully resolving the two targets using the periodogram, MUSIC, and ESPRIT algorithms for $\mathrm{SNR}=-10\mathrm{dB}$ and $\mathrm{SNR}=10\mathrm{dB}$ are shown in Figs. 17(a) and 17(b), respectively, based on $200$ Monte Carlo experiments.
The results in Figs. 17(a) and 17(b) show the angular resolution of the periodogram algorithm is $7^{\circ}$ for both $\mathrm{SNR}=-10\mathrm{dB}$ and $\mathrm{SNR}=10\mathrm{dB}$, consistent with the theoretical resolution in (41). The resolution of the MUSIC and ESPRIT algorithms depends on the SNR. For lower SNR, MUSIC and ESPRIT exhibit similar resolution, approximately $4^{\circ}$. For higher SNR, ESPRIT achieves superior resolution around $0.6^{\circ}$, whereas MUSIC provides resolution around $1^{\circ}$. Overall, as expected, super-resolution algorithms significantly enhance parameter estimation resolution compared to the periodogram, particularly at high SNR levels.
@@
IV-F Lessons Learned
In far-field scenarios, the full decoupling of the three-domain signal processing and the linear array-manifold-like structure in each domain provide significant flexibility for target sensing. Based on the four methods depicted in Fig. 6 and different radar sensing algorithms, various trade-offs between complexity, resolution, and accuracy can be achieved. In particular, these methods can accomplish target sensing under very low SNR conditions, which exacerbates the disparity between C&S coverage. Specifically, communication typically relies on the LoS and multipaths between the BS and users, whereas sensing depends on the two-hop path from the ISAC-TX to the target and then to the Sen-RX. Consequently, the communication SNR is generally higher than that for sensing. However, there are also significant differences in the SNR requirements for C&S signal processing. This discrepancy in both signal SNR and SNR requirements between C&S leads to differences in their coverage ranges, which warrants further in-depth investigation.
For spatial sensing, super-resolution sensing algorithms are not directly applicable to ISAC systems equipped with analog or hybrid beamforming-based MIMO arrays. Nevertheless, beamforming optimization holds promise for facilitating narrower beams than the Rayleigh limit, thereby enhancing spatial sensing resolution. For target detection, non-super-resolution algorithms can apply constant false alarm rate (CFAR) [140] detection to target sensing periodograms to accurately estimate the number of targets. In contrast, subspace-based super-resolution algorithms divide the signal and noise subspaces based on the magnitude of the eigenvalues of the covariance matrix, which may lead to higher missed detection or false alarm rates. Target detection methods for super-resolution sensing algorithms still require further exploration. Finally, implementing super-resolution sensing algorithms on FPGA chips for efficient computation remains challenging and requires more research effort.
@@ V Near-field MIMO-OFDM ISAC
In this section, we extend the results in Section IV to the case where the targets are located in the near-field region of the Sen-RX. While the main signal processing strategies outlined in Fig. 6 are still valid in the near-field, the spatial domain estimation must be extended from AoA estimation only in the far-field to joint angle-range estimation in the near-field by exploiting the characteristics of spherical wavefronts. Therefore, this section focuses mainly on spatial domain estimation.
In the near-field region, propagation is described by spherical wavefronts, which produce a nonlinear variation in phase from element to element of the array. By leveraging the distance information embedded in the spherical wavefront, near-field beamforming can focus energy precisely at a designated location. Therefore, the energy can be concentrated in both the angle and distance dimensions, producing a narrow beam width (BW) and beam depth (BD), respectively [141], as shown in Fig. 18. In wideband scenarios, near-field propagation causes the beams at different frequencies to point to different physical locations, an effect referred to as near-field beam splitting. While this effect may result in significant loss of array gain since it is difficult to achieve beam alignment across a wide frequency band, one can control the angular coverage of the beams at different frequencies in a way that enables rapid CSI acquisition [142].
Figure 18: Illustration of beam characteristics in near-field ISAC.
@@ V-A Near-field Sensing Signal Model
As shown in Fig. 19, we assume that the Sen-RX is equipped with an extremely large-scale ULA, which is placed along the $y$-axis with $\varepsilon_{r}=0,$ and $\varepsilon_{m}=m-(M_{r}+1)/2$ in (14). The corresponding steering matrix is denoted as $\mathbf{A}_{\rm{R}}(r,\theta)=[\mathbf{a}_{\mathrm{R}}(r_{1},\theta_{1}),\cdots,\mathbf{a}_{\mathrm{R}}(r_{K},\theta_{K})]$. Based on the model in (22), the $M_{r}\times Q_{N,P}$ received signal is given by
@@
where $\mathbf{S}=[\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{K}]^{T}\in\mathbb{C}^{K\times Q_{N,P}}$ denotes the equivalent reflected target signals described in (23). Note that compared with the far-field model in (22), the array response vector for near-field sensing is a function of both the target angle and distance, hence allowing simultaneous estimation of both parameters independent of delay-based distance estimation. This will not only improve the distance estimation accuracy (especially if the distance estimation is performed jointly in the AoA and delay domains), but it can also alleviate the synchronization errors that typically arise in traditional delay-based distance estimation techniques. In order to estimate the angle and distance parameters based on spatial-domain processing, several methods are introduced below that either ignore the near-field model and directly apply far-field algorithms or extend far-field algorithms to the near-field case.
V-B Directly Applying Far-field Algorithms to Near-field
In this section, we illustrate the performance of directly applying the far-field algorithms discussed in Section IV-B to near-field scenarios while ignoring the spherical wave propagation. The result for the standard Periodogram is shown in Fig. 20, where $K=2$ targets are considered. We see that only the target with $r=150m$ has an accurate AoA estimate based on (45). For the nearer target at $r=30m$, multiple peaks exist around the ground-truth AoA, and it is due to the mismatch between the near-field channel and the assumed far-field model. This is known as the so-called energy spread effect, which illustrates the necessity of developing dedicated near-field target localization algorithms.
@@ V-B Directly Applying Far-field Algorithms to Near-field
V-C Extending Far-field Algorithms to Near-field
@@ V-C1 Near-field Beam Focusing
For near-field models, the distance and angle parameters are closely coupled in the near-field array response vector. In addition, the non-linear phase variation from antenna to antenna in near-field models also means that the IFFT is no longer directly applicable. Therefore, in this section we introduce 2D beamforming in both the angle and distance domains. Based on (74), the spatial domain sample covariance matrix of $\bf X$ is given by
@@
The peaks of $P_{\mathrm{2D-BF}}$} in the 2D angle-distance plane can be used to estimate the target angle and distance parameters.
@@ V-C2 Near-field MUSIC
Similar to (47), by performing the EVD of the covariance matrix ${\bf R}_{X}$ and obtaining the noise subspace ${\bf E}_{n}$, the 2D MUSIC spectrum can be expressed as
@@ V-C4 Simulations and discussions
Figs. 21(a) and 21(b) show the 2D spectra for near-field beam focusing and MUSIC respectively. The number of antennas is $M_{r}=256$, and the locations of the two targets are $(r_{1},\theta_{1})=(5\mathrm{m},10^{\circ})$ and $(r_{2},\theta_{2})=(10\mathrm{m},20^{\circ})$, respectively. It is observed that both beamforming and MUSIC can localize the closer target with higher accuracy than the distant target, primarily due to the limited accuracy with respect to distance.
@@
Figure 21: Near-field angle-distance estimation.
We see that several issues arise that complicate parameter estimation in near-field scenarios. Besides the fact that some algorithms can no longer be directly applied, we see that 2D rather than 1D searches are necessary to jointly estimate the AoAs and ranges, which adds complexity. The accuracy of the distance estimate is considerably less reliable than the AoA estimate, especially as the distance increases to the point where the far-field model holds [143]. This is due to the decreasing sensitivity of the model with distance in such scenarios, and is a factor that must be taken into account when jointly estimating the distance by including the delay domain.
To address these challenges, one feasible approach is to decouple estimation of the angle and distance parameters to reduce the complexity of implementing the parameter optimization. The low-complexity near-field sensing algorithms presented next are based on this idea.
@@ V-D Low-complexity Near-Field Estimation Algorithms
Recently, several algorithms have been proposed to improve the accuracy and resolution of near-field sensing with low computational complexity by efficiently decoupling the distance and angle parameters. Many such methods exploit the approximate Fresnel array response model in (16) based on a second-order Taylor approximation, where the parameters to be estimated are taken to be ${\omega_{k}}=-\frac{{2\pi{d}\sin{\theta_{k}}}}{\lambda}$ and ${\psi_{k}}=\frac{{\pi d^{2}{{\cos}^{2}}{\theta_{k}}}}{{\lambda{r_{k}}}}$ instead of $\theta_{k}$ and $r_{k}$.
@@ V-D1 Second-order cumulant (SoC)
Note that in (16), $\omega_{k}$ only depends on $\theta_{k}$, while $\psi_{k}$ depends on both $\theta_{k}$ and $r_{k}$. Therefore, a dedicated SoC can be constructed to transform the multi-dimensional search into two single-parameter estimation problems. Based on (74) and (16), the noise-free correlation sequence between the $(am+b)$th and $m$th antenna elements can be written as [144]
@@
where ${\bf{A}}_{R,m}$} is the $m$th row of ${\bf{A}}_{R}$}. Approximation $(a)$ holds by assuming ${\bf{s}}_{p}^{H}{{\bf{s}}_{q}}\approx 0,p\neq q$. $J=\frac{M_{r}-1}{2}$ and $p_{k}=\frac{1}{Q_{N,P}}{\bf{S}}(k,:){{\bf{S}}^{H}}(k,:)$ denotes the equivalent transmitted signal. Furthermore, by properly choosing $a$ and $b$, $\psi_{k}$ can be eliminated from the resulting expression, and estimation of the distance and angle parameters can be decoupled. For instance, substituting (a,b)=(1,0)(a,b)=(-1,0) and $(a,b)=(1,1)$ into (V-D1), we obtain the following correlation sequences
@@ V-D3 FFT-enhanced low-complexity and super-resolution algorithm
In order to further reduce the complexity, the authors in [149] proposed an efficient algorithm to narrow the search region for 2D angle-distance estimation in (77). The main idea of this algorithm is to first eliminate those regions that do not contain any targets. Afterwards, 2D MUSIC is performed within a region of reduced size, effectively balancing computational complexity and super-resolution. This algorithm involves three steps:
FFT-enhanced angle cluster determination: Applying an $S$-point FFT and IFFT to the columns and rows of sample covariance matrix ${\bf R}_{X}$ in (75), respectively, the output can be used to eliminate angular regions where no targets exist. The resulting spectrum can be expressed as ${{\bf{p}}}=\text{diag}({\bf{W}}{{\bf{R}}_{X}}{{\bf{W}}^{-1}}),$ where $\bf W$ is the DFT matrix, and diag(A) extracts the diagonal elements of $\bf A$ to form a vector. The $n$th angle cluster is represented by the minimum and maximum angle value within this cluster, denoted as ${\bm{\alpha}}_{n}=[{\underline{\alpha}_{n}},{{\overline{\alpha}}_{n}}]$ for $n=1,\cdots,L$, where $L$ is the number of clusters.
Distance cluster determination: For each angle cluster ${\bm{\alpha}}_{n}=[{\underline{\alpha}_{n}},{{\overline{\alpha}}_{n}}]$, 1D beamforming is performed over the distance domain for $\underline{\alpha_{n}}$ and ${\overline{\alpha}}_{n}$},
@@
By setting a threshold $\Gamma$, the distance cluster ${{\bm{\underline{\beta}}}}_{n}$} can be determined by the condition $P_{low}(r)<\Gamma$, and ${{\bm{\bar{\beta}}}}_{n}$} can be obtained similarly. Therefore, the distance cluster is expressed as ${\bm{\beta}}_{n}={{{\bm{\underline{\beta}}}}_{n}}\cup{{{\bm{\bar{\beta}}}}_{n}}.$
Super-resolution 2D sensing: Finally, for each cluster $n$, 2D MUSIC is performed for $\theta\in{\bm{\alpha}}_{n}$ and $r\in{\bm{\beta}}_{n}$. By clustering the angle and distance dimensions, the number of grid points for the 2D search and the overall complexity is significantly reduced.